A body of mass 50 kg is moving with a constant velocity of 5m/s on a horizontal, frictionless surface in free space.Calculate the force acting on the body.Account for the obtained value of force.

Grade:9

1 Answers

Vikas TU
7 years ago
F = m*a
F = m*dv/dt
dv =  (F/m) dt
integerating both sides,
v = (F/m)t
F = 5m/t = > 5 * 50 / t = 2150/t
put t at which u want to get the force value.
.
And here the contradiction is but
the velocity is constant therefore a = 0
hence extrenal force = 0.